OpenStudy (brittany):

rationl or irrational √.25

OpenStudy (anonymous):

sqrt(25/100) is rational

OpenStudy (anonymous):

is rational since 25<100 and 1<4

OpenStudy (amistre64):

if you can get rid of the sqrt part, it can be ratioed

OpenStudy (anonymous):

5/10=.5

OpenStudy (anonymous):

it is rational since it is a perfect square. that is, it is the square of a rational number. it is the square of \[\frac{1}{2}\]

OpenStudy (anonymous):

since \[(\frac{1}{2})^2=\frac{1}{4}=.25\]